A race car accelerates uniformly from 18.5 m/s to 46.1m/s in 2.47 seconds. Determine the acceleration of the car and the distance traveled.
 

Grade:12

1 Answers

Trap
2 years ago

Here, initial velocity is u=18.5m/s and final velocity is v=46.1m/s 

Time taken t=2.47s
If a be the acceleration of the car.
Using v=u+at,
46.1=18.5+a(2.47)
 a=11.17m/s2
If S be the distance traveled by car.
Using formula v2u2=2aS,
(46.1)2(18.5)2=2(11.17)S
 S=79.8m